# Kummer-Artin-Schreier-Witt Theory

### v2

@@ 4.2.1 Approximation of the Power Series $E_{s,p}$
    If $r>1$: This case arises when $n$ is not divisible by $p$. Then $\nu(m_{k}!)=0$, so the valuation becomes $1+\sum_{j=1}^{r}m_{j}\nu(a_{i_{j}})$.
Assume there exists an index $m>\frac{l}{p}$ such that $\nu(a_{m})<\frac{1}{p-1}$. Then $a_{m}^{p}$ contributes to $c_{pm}$. For $\nu(c_{pm})\geq\frac{p}{p-1}$, there must be distinct indices $i_{1},\ldots,i_{s}<pm$ such that $\sum_{j=1}^{r^{\prime}}m_{j}i_{j}=pm$, where $\sum_{j=1}^{r^{\prime}}m_{j}=p$, which implies that
$1+\sum_{j=1}^{r^{\prime}}m_{j}\nu(a_{i_{j}})\leq p\nu(a_{m}),$ (25)
which is impossible if $\nu(a_{m})\leq\frac{1}{p}$. Thus, we conclude that $\nu(a_{m})>\frac{1}{p}$ for $m>\frac{l}{p}$.
Additionally, due to the assumption $\nu(a_{m})<\frac{1}{p-1}$, we have the relation
$p\nu(a_{m})-1-\nu(a_{m})=(p-1)\nu(a_{m})-1<1-1=0,$
which implies $p\nu(a_{m})-1<\nu(a_{m})$.
Furthermore, for $\sum_{j=1}^{r^{\prime}}m_{j}i_{j}=pm>l$, at least one $i_{s}>\frac{l}{p}$ for some $s$. Therefore, $\nu(a_{i_{s}})>\frac{1}{p}$. Finally, as the lower bound of the valuations is $B$, it follows from (25) that
$\nu(a_{i_{s}})\leq p\nu(a_{m})-1-(p-1)B<\nu(a_{m})-(p-1)B.$
This leads to coefficients with progressively smaller valuations until reaching a value not exceeding $\frac{1}{p}$. Therefore, we can conclude that $\nu(a_{m})\geq\frac{1}{p-1}$ for all $m>\frac{l}{p}$, completing the proof. ∎
@@
Proposition 4.4 extends to the case $f\in R\llbracket X_{0},\ldots,X_{i-1}\rrbracket$, where $i>1$.
@@
We may view $f$ as a series in $R\llbracket X_{0},\ldots,X_{i-2}\rrbracket\llbracket X_{i-1}\rrbracket=:R_{i-1}\llbracket X_{i-1}\rrbracket$, where $R_{i-1}$ is equipped with the Gauss valuation. Applying Proposition 4.4 to $f$ in terms of $X_{i-1}$ then allows us to assume that
$f=1+\sum_{j=1}^{n}a_{j}X_{i-1}^{j}+O\left(p^{1/(p-1)}\right),$
where $a_{j}\in R_{i-1}$, and $\nu(a_{j})<1/(p-1)$ for $j=1,\ldots,n$. Furthermore, we have
$f^{p}=(1+\sum_{j=1}^{n}a_{j}X_{i-1}^{j})^{p}+O\left(p^{p/(p-1)}\right).$ (26)
Therefore, $f_{i-1}^{p}:=(1+\sum_{j=1}^{n}a_{j}X_{i-1}^{j})^{p}$ has finitely many terms with valuations smaller than $p/(p-1)$. We then regard $f_{i-1}$ as a series in $R\llbracket X_{0},\ldots,X_{i-3},X_{i-1}\rrbracket\llbracket X_{i-2}\rrbracket=:R_{i-2}\llbracket X_{i-2}\rrbracket$. Applying Proposition 4.4 to $f_{i-1}$ in terms of $X_{i-2}$, we find that
$f_{i-1}=1+\sum_{j=1}^{m}b_{j}X_{i-2}^{j}+O\left(p^{1/(p-1)}\right),$
where $b_{j}\in R\llbracket X_{0},\ldots,X_{i-3}\rrbracket[X_{i-1}]$. Repeating this process $i-2$ times completes the proof. ∎
@@
$E_{s,p}(p(Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))$ contains only finitely many terms with coefficients of valuation strictly smaller than $p/(p-1)$.
@@
The first factor meets the condition because it is a polynomial, and the second factor does so by Lemma 2.4 and the fact that $Y_{s}^{\prime}$ satisfies the conditions of Proposition 4.2 by the induction hypothesis. The proposition then follows. ∎
@@
The result follows from Proposition 4.3, Proposition 4.5, and Proposition 4.6. ∎
@@ 4.2.2 Computing the Minimal Polynomial
$Y_{s-1}^{\prime}=(-1)^{p}\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+F_{s}(Y_{0},\ldots,Y_{s-1}),$
@@
The first two expressions are series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-2}\rrbracket$, and the last is a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\allowbreak\ldots,X_{s-2}\rrbracket$. Additionally, for $j=0,\ldots,s-1$, we have $X_{j}^{\prime}=Y_{j}^{\prime}$, and $X_{j}$ can be expressed as a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{j-1}}]\left[Y_{0},\ldots,Y_{j},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{j-1}+\lambda Y_{j-1}}\right]$ by induction. Thus, we obtain
@@
as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-1}\rrbracket$. This forces the side to be a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\big{[}Y_{0},\ldots\allowbreak\ldots,Y_{s-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-2}+\lambda Y_{s-2}}\big{]}$. Furthermore, the left-hand side is a degree-$p$ expression in $Y_{s-1}$, so it is the minimal polynomial of $Y_{s-1}$ over $\mathbb{Q}[\zeta_{p^{s}}](Y_{0},\ldots,Y_{s-2},X_{s-1})$. Finally, the claim that $X_{s-1}$ is an element of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\left[Y_{0},\ldots,Y_{j},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-1}+\lambda Y_{s-2}}\right]$ follows immediately. ∎
@@
Since $Y_{s}^{\prime}=X_{s}^{\prime}$, there exists a power series $H_{s-1}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\ldots,X_{s-1}\rrbracket$, which is also a unit, such that
$\begin{split}-uY_{s-1}^{p}+Y_{s-1}+F_{s}&=X_{s-1}H_{s-1},\\ H_{s-1}^{-1}(-uY_{s-1}^{p}+Y_{s-1}+F_{s})(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})&=X_{s-1}(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}),\end{split}$
where $u=\frac{\lambda^{p-1}}{p}(-1)^{p}$. By Proposition 4.11, we can regard both $X_{s-1}$ and $H_{s-1}^{-1}$ as elements of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-1}\rrbracket$. Thus, $H_{s-1}^{-1}\equiv 1\pmod{(\zeta_{p^{s}}-1)}$. Consequently, $-uY_{s-1}^{p}+Y_{s-1}+F_{s}$ is equivalent to the left-hand side of (28) modulo $(\zeta_{p^{s}}-1)$.
The result then follows from the expansion
@@
and the fact that $E_{s}(Y_{0},\ldots,Y_{s-2})$ reduces to $1$ modulo $(\zeta_{p^{s}}-1)$. ∎
That concludes the proof of Proposition 4.2. ∎
@@
For $s>1$, recall that the special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the extension of $\mathbb{F}_{p}[x_{0},\ldots,x_{s-1}]$ associated with the Witt vector $(x_{0},\ldots,x_{s-1})$, which can be defined as a system of Artin-Schreier extensions as in (8). By induction, the last of these extensions is given by
@@
for some $b_{s-1}\in\mathbb{F}_{p}[y_{0},\ldots,y_{s-2},x_{s-1}]$. On the other hand, by (27), we have
@@
where $\overline{H}_{s}$ denotes the reduction of $H_{s}$ modulo $(\zeta_{p^{s}}-1)$. The result follows immediately. ∎
@@
When $p=2$, $\overline{H}_{2}=y_{0}^{4}+y_{0}^{8}=0\wp(y_{0})^{4}$ if we choose $G_{2}(Y_{0})=1+2\zeta_{4}Y_{0}-52Y_{0}^{4}$. This is the lowest degree polynomial for $G_{2}$ that we can select.
@@ 5.1 The Case $s=1$
The covering $\mathcal{W}_{1}$ is defined by the diagram in Figure 1, showing the relationships between $\mathcal{W}_{1}$, $\mathcal{V}_{1}$, and their function fields.
[…]
@@
That is, the coordinate ring of $\mathcal{W}_{1}$ is the integral closure of that of $\mathcal{V}_{1}$ in $L_{1}$. Recall also that applying the change of variables $Z_{0}=1+\lambda Y_{0}$ to (29) yields
@@
Thus, $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$ has good reduction $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]$. Additionally, from (29) and Kummer theory, the cover $\mathcal{Y}_{1}\rightarrow\mathbb{P}^{1}_{R}$ branches generically at $X_{0}=\infty$ and $X_{0}=\frac{1}{p\lambda}$. Therefore, $\mathcal{W}_{1}\rightarrow\mathcal{V}_{1}$ is étale with Galois group $\mathbb{Z}/p$, and its special fiber is isomorphic to the Artin-Schreier extension $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]\rightarrow\operatorname{Spec}\mathbb{F}_{p}[x_{0}]$ given by $-y_{0}^{p}+y_{0}=x_{0}$.
5.2 When $s>1$
Consider Figure 2, which extends Figure 1 to show the step-by-step construction of $\mathcal{W}_{s}$.
[…]

### v3

@@ 4.2.1 Approximation of the Power Series $E_{s,p}$
    If $r>1$: Then $\nu(m_{k}!)=0$, so the valuation becomes $1+\sum_{j=1}^{r}m_{j}\nu(a_{i_{j}})$.
Set
$l:=\max\{n\mid\nu(c_{n})<\tfrac{p}{p-1}\},$
with the convention l:=1l:=-1 if the set is empty. Assume there exists an index $m>\frac{l}{p}$ such that $\nu(a_{m})<\frac{1}{p-1}$. Then $a_{m}^{p}$ contributes to $c_{pm}$, and
$\nu(a_{m}^{p})=p\nu(a_{m})<\frac{p}{p-1}.$
Since $pm>l$, we have $\nu(c_{pm})\geq\frac{p}{p-1}>p\nu(a_{m})$. By the non-archimedean property, if $\nu\!\left(\sum_{\alpha}t_{\alpha}\right)>\min_{\alpha}\nu(t_{\alpha})$ then the minimum $\min_{\alpha}\nu(t_{\alpha})$ is attained by at least two summands. Applying this to the sum of multinomial summands defining $c_{pm}$, we deduce that there exists a summand with $r^{\prime}>1$ (hence with multinomial coefficient of valuation $1$) and distinct indices $i_{1},\ldots,i_{r^{\prime}}\leq pm$ such that $\sum_{j=1}^{r^{\prime}}m_{j}i_{j}=pm$, $\sum_{j=1}^{r^{\prime}}m_{j}=p$, and
$1+\sum_{j=1}^{r^{\prime}}m_{j}\nu(a_{i_{j}})\;\leq\;p\nu(a_{m}).$ (27)
In particular, (27) is impossible if $\nu(a_{m})\leq\frac{1}{p}$, because the left-hand side is >1>1 (at least one index satisfies $i_{j}>0$ since $\sum m_{j}i_{j}=pm>0$, so one term has $\nu(a_{i_{j}})>0$) whereas the right-hand side is $\leq 1$. Hence $\nu(a_{m})>\frac{1}{p}$. Also, since $\nu(a_{m})<\frac{1}{p-1}$, we have
$p\nu(a_{m})-1<\nu(a_{m}).$
From $\sum_{j=1}^{r^{\prime}}m_{j}i_{j}=pm>l$, at least one index satisfies $i_{s}>\frac{l}{p}$. We claim $\nu(a_{i_{s}})<\frac{1}{p-1}$: otherwise
$1+\sum_{j=1}^{r^{\prime}}m_{j}\nu(a_{i_{j}})\geq 1+\nu(a_{i_{s}})\geq\frac{p}{p-1},$
contradicting (27) together with $p\nu(a_{m})<\frac{p}{p-1}$. By the same argument as above applied to $i_{s}$, we also have $\nu(a_{i_{s}})>\frac{1}{p}$. Finally, (27) gives
$\nu(a_{i_{s}})\leq p\nu(a_{m})-1-\sum_{j\neq s}m_{j}\nu(a_{i_{j}})-(m_{s}-1)\nu(a_{i_{s}})\leq p\nu(a_{m})-1<\nu(a_{m}).$
Thus, starting from $m_{0}:=m$, we recursively obtain indices $m_{t+1}>\frac{l}{p}$ such that
$\frac{1}{p}<\nu(a_{m_{t+1}})<\nu(a_{m_{t}})<\frac{1}{p-1}.$
This gives an infinite strictly decreasing sequence in the discrete value group $\nu(K^{\times})$, impossible. Hence no such $m$ exists, so $\nu(a_{m})\geq\frac{1}{p-1}$ for all $m>\frac{l}{p}$, as required. ∎
@@
Proposition 4.4 extends to the multivariable case as follows. Let $i>1$ and write
$f=\sum_{n\in\mathbb{Z}_{\geq 0}^{i}}a_{n}X^{n}\in R\llbracket X_{0},\ldots,X_{i-1}\rrbracket,\qquad X^{n}:=X_{0}^{n_{0}}\cdots X_{i-1}^{n_{i-1}}.$
Assume $a_{\mathbf{0}}=1$, $\nu(a_{n})>0$ for $n\neq\mathbf{0}$, and that the set $\{\nu(a_{n})\mid n\neq\mathbf{0}\}$ is bounded below by some $B>0$. If $f^{p}$ has only finitely many monomials whose coefficients have valuation $\frac{p}{p-1}$, then $f$ has only finitely many monomials whose coefficients have valuation $\frac{1}{p-1}$.
@@
For each $r\in\{0,\ldots,i-1\}$, write
$f=\sum_{j\geq 0}A_{r,j}\,X_{r}^{j},\qquad A_{r,j}\in R\llbracket X_{0},\ldots,\widehat{X_{r}},\ldots,X_{i-1}\rrbracket.$
Equip the coefficient ring with the Gauss valuation (still denoted $\nu$)
$\nu\!\left(\sum b_{m}X^{m}\right):=\inf_{m}\nu(b_{m}).$
Then $A_{r,0}$ is a unit, $\nu(A_{r,j})>0$ for $j>0$, and these valuations are bounded below by $B>0$.
Write also
$f^{p}=\sum_{j\geq 0}C_{r,j}X_{r}^{j}.$
If infinitely many $j$ satisfied $\nu(C_{r,j})<\tfrac{p}{p-1}$, then for each such $j$ we can choose a monomial term of $C_{r,j}$ whose coefficient attains this valuation (the value group is discrete). This yields infinitely many monomials of $f^{p}$ with coefficient valuation $\tfrac{p}{p-1}$, contradicting the hypothesis. Hence, for each fixed $r$, only finitely many $j$ satisfy $\nu(C_{r,j})<\tfrac{p}{p-1}$. Applying Proposition 4.4 to the one-variable series $\sum_{j\geq 0}A_{r,j}X_{r}^{j}$ (over the Gauss-valued coefficient ring) gives that
$J_{r}:=\{j\geq 0\mid\nu(A_{r,j})<\tfrac{1}{p-1}\}$
is finite for each $r$.
Now let $a_{n}X^{n}$ be a monomial of $f$ with $\nu(a_{n})<\tfrac{1}{p-1}$. For each $r$, the coefficient $a_{n}$ appears in $A_{r,n_{r}}$, so
$\nu(A_{r,n_{r}})\leq\nu(a_{n})<\tfrac{1}{p-1},$
hence $n_{r}\in J_{r}$. Therefore
$n\in J_{0}\times\cdots\times J_{i-1},$
and the right-hand side is finite. So $f$ has only finitely many monomials whose coefficients have valuation $\tfrac{1}{p-1}$. ∎
@@
$E_{s,p}(p(Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))$ contains only finitely many terms with coefficients of valuation strictly smaller than p/(p1)p/(p-1).
@@
The second equality is the defining relation $Z_{s-2}=E_{s-1,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime}))$ together with the change of variables (26) for i=s1i=s-1. The first factor meets the finiteness condition because it is a polynomial. For the second factor, Lemma 2.5 shows that
$G_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket(\zeta_{p^{s}}-1)Y_{0}^{\prime},\ldots,(\zeta_{p^{2}}-1)Y_{s-2}^{\prime}\rrbracket.$
Since each $\zeta_{p^{m}}-1$ has positive valuation and $Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime}$ have coefficients in $R$ (by the induction hypothesis in Proposition 4.2), only finitely many monomials in this expansion can have coefficient valuation $\frac{p}{p-1}$. The proposition follows. ∎
@@
Set $f:=E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))\in R\llbracket Y_{0},\ldots,Y_{s-2}\rrbracket$. By Proposition 4.3, $f$ satisfies the hypotheses of Proposition 4.5. Moreover, $f^{p}=E_{s,p}(p(Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))$, and Proposition 4.6 shows that $f^{p}$ has only finitely many terms with coefficient valuation $\frac{p}{p-1}$. Applying Proposition 4.5 to $f$ yields the claim. ∎
@@ 4.2.2 Computing the Minimal Polynomial
$Y_{s-1}^{\prime}=(-1)^{p-1}\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+F_{s}(Y_{0},\ldots,Y_{s-1}),$
@@
The first two expressions are series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-2}\rrbracket$, and the last is a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\allowbreak\ldots,X_{s-2}\rrbracket$. Additionally, for $j=0,\ldots,s-1$, we have $X_{j}^{\prime}=Y_{j}^{\prime}$, and $X_{j}$ can be expressed as a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{j+1}}]\left[Y_{0},\ldots,Y_{j},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{j-1}+\lambda Y_{j-1}}\right]$ by induction. Thus, we obtain
@@
as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket Y_{0},\ldots,Y_{s-1}\rrbracket$. This forces the left-hand side to be a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\big[Y_{0},\ldots,Y_{s-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-2}+\lambda Y_{s-2}}\big]$. Thus the left-hand side can be written as
$\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+H_{s}(Y_{0},\ldots,Y_{s-1})$
for some $H_{s}$ in that ring. Furthermore, the left-hand side is a degree-$p$ expression in $Y_{s-1}$, and the Kummer equation defines a degree-$p$ extension over $\mathbb{Q}[\zeta_{p^{s}}](Y_{0},\ldots,Y_{s-2},X_{s-1})$, so this polynomial is the minimal polynomial of $Y_{s-1}$. Finally, for j=s1j=s-1 the defining equation (28) explicitly solves for $X_{s-1}$ by dividing by $(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})$, which is one of the inverted factors in the stated ring; the cases $j\leq s-2$ follow by induction as stated. ∎
@@
Set $\pi:=\zeta_{p^{s}}-1$ and $u:=\frac{\lambda^{p-1}}{p}(-1)^{p}$. By Proposition 4.10 and the identity $Y_{s-1}^{\prime}=X_{s-1}^{\prime}$, we have
$uY_{s-1}^{p}+Y_{s-1}+F_{s}=X_{s-1}^{\prime}.$
By Proposition 3.5, $X_{s-1}^{\prime}\equiv X_{s-1}\pmod{\pi}$, hence
$uy_{s-1}^{p}+y_{s-1}+\overline{F}_{s}=x_{s-1}.$ (30)
On the other hand, reducing (28) modulo $\pi$ and using
$E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}\equiv 1\pmod{\pi}$,
we obtain
$uy_{s-1}^{p}+y_{s-1}+\overline{H}_{s}=x_{s-1}.$ (31)
Comparing (30) and (31) gives $\overline{F}_{s}=\overline{H}_{s}$.
It remains to show $\overline{H}_{s}\in\mathbb{F}_{p}[y_{0},\ldots,y_{s-2}]$. From (29),
@@
where the omitted mixed terms have valuation $\geq p/(p-1)$. After dividing by $p\lambda$ and reducing modulo $\pi$, these omitted terms vanish, and $E_{s}(Y_{0},\ldots,Y_{s-2})\equiv 1\pmod{\pi}$ by Definition 4.9. Therefore $\overline{H}_{s}$ is a polynomial in $y_{0},\ldots,y_{s-2}$, so the same holds for $\overline{F}_{s}$. ∎
This completes the proof of Proposition 4.2.
@@
For $s>1$, recall that the special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is birational to the extension of $\mathbb{F}_{p}[x_{0},\ldots,x_{s-1}]$ associated with the Witt vector $(x_{0},\ldots,x_{s-1})$, which can be defined as a system of Artin–Schreier extensions as in (9). By induction, the last of these extensions is given by
@@
where $b_{j-1}\in\mathbb{F}_{p}[y_{0},\ldots,y_{j-2}]$ for $j=2,\ldots,s-1$ by the induction hypothesis, and a priori we may take $b_{s-1}\in\mathbb{F}_{p}[y_{0},\ldots,y_{s-2},x_{s-1}]$. On the other hand, by (28), we have
@@
where $\overline{H}_{s}$ denotes the reduction of $H_{s}$ modulo $(\zeta_{p^{s}}-1)$. Comparing with the previous displayed equation (and cancelling the common terms $y_{s-1}-y_{s-1}^{p}$ and $x_{s-1}$), we obtain
$\overline{H}_{s}=f_{s}(y_{0},y_{1}+b_{1},\ldots,y_{s-2}+b_{s-2})+\wp(b_{s-1}).$
In particular, $\wp(b_{s-1})\in\mathbb{F}_{p}[y_{0},\ldots,y_{s-2}]$, so we may (and do) choose $b_{s-1}\in\mathbb{F}_{p}[y_{0},\ldots,y_{s-2}]$. ∎
@@
When $p=2$, $\overline{H}_{2}=y_{0}^{2}-y_{0}^{3}+y_{0}^{4}-y_{0}^{8}=f_{2}(y_{0})+\wp(y_{0}^{4})$ if we choose $G_{2}(X_{0})=1+2\zeta_{4}X_{0}-52X_{0}^{4}$. This is the lowest degree polynomial for $G_{2}$ that we can select.
@@ 5.1 The case $s=1$
The cover $\mathcal{W}_{1}$ is defined by the diagram in Figure 1, which records the relationships between $\mathcal{W}_{1}$, $\mathcal{V}_{1}$, and their function fields.
$\mathcal{W}_{1}$$\mathcal{Y}_{1}$$\operatorname{Spec}L_{1}$$\mathcal{V}_{1}$$\mathbb{P}^{1}_{R}$$\operatorname{Spec}K(X_{0})$}
@@
That is, the coordinate ring of $\mathcal{W}_{1}$ is the integral closure of that of $\mathcal{V}_{1}$ in $L_{1}$. Substituting $Z_{0}=1+\lambda Y_{0}$ into (32) yields
@@
Thus, $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$ has good reduction, with special fiber $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]$. Moreover, by (32) and Kummer theory, the cover $\mathcal{Y}_{1}\rightarrow\mathbb{P}^{1}_{R}$ is generically branched at $X_{0}=\infty$ and $X_{0}=-\frac{1}{p\lambda}$. Therefore, $\mathcal{W}_{1}\rightarrow\mathcal{V}_{1}$ is étale with Galois group $\mathbb{Z}/p$, and its special fiber is isomorphic to the Artin–Schreier extension $\operatorname{Spec}\mathbb{F}_{p}[y_{0}]\rightarrow\operatorname{Spec}\mathbb{F}_{p}[x_{0}]$ given by $y_{0}^{p}+y_{0}=x_{0}$.
5.2 The case $s>1$
Consider Figure 2, which extends Figure 1 and shows the step-by-step construction of $\mathcal{W}_{s}$.
$\mathcal{W}_{s}$$\mathcal{Y}_{s}$$\operatorname{Spec}L_{s}$$\mathcal{W}_{s-1}$$\mathcal{Y}_{s-1}$$\operatorname{Spec}L_{s-1}$$\vdots$$\vdots$$\vdots$$\mathcal{W}_{1}$$\mathcal{Y}_{1}$$\operatorname{Spec}L_{1}$$\mathcal{V}_{s}$$\mathbb{P}^{s}_{R}$$\operatorname{Spec}K(X_{0},X_{1},\ldots,X_{s-1})$}